OpenStudy (anonymous):

how do you do this? sin(x/2)=1-cos x

OpenStudy (badhi):

use the property- $$\cos(2x)=1-2\sin^2x$$

OpenStudy (aravindg):

you mean \[\large \sin^2 (\dfrac{x}{2})?\]

OpenStudy (badhi):

that is for $$\cos(x)$$

OpenStudy (badhi):

I have mentioned it for $$\cos(2x)$$